OpenStudy (anonymous):

Describe the differences between the graph of y = –3(x + 7)2 – 10 and the standard position graph of y = x2.

OpenStudy (anonymous):

*y=-3(x+7^2-10 , *y=x^2

OpenStudy (anonymous):

y=-3(x+7)^2-10 y=x^2

OpenStudy (anonymous):

is that your equation

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i tried to draw but the drawin thing suxs.... the first equation as you can see is like this y=-ax^2-10 so here a is negative so as x goes along the x axis from zero , y becomes a bigger negative number, also when  y=x^2